OpenStudy (anonymous):

5(-3x - 2) - (x - 3) = -4(4x + 5) + 13

OpenStudy (anonymous):

5(-3x - 2) - (x - 3) = -4(4x + 5) + 13 -15x - 10 - x + 3 = -16x - 20 + 13 -16x - 7 = -16x - 7 0 = 0